Suppose J is between H and K. Use the Segment Addition Postulate to solve for x. Then find the length of each segment if HJ=1/3x + 4 , JK = 2x + 2/3, KH = 2 2/3x + 1.
Answer by Nate(3500)

H ---- J --------- K
HJ = x/3 + 4
JK = 2x + 2/3
KH = 8x/3 + 1
x/3 + 4 + 2x + 2/3 = 8x/3 + 1
x + 12 + 6x + 2 = 8x + 3
7x + 14 = 8x + 3
11 = x
HJ = 11/3 + 4 = 11/3 + 12/3 = 23/3
JK = 2(11) + 2/3 = 66/3 + 2/3 = 68/3
HK = 8(11)/3 + 3/3 = 88/3 + 3/3 = 91/3